OpenStudy (kamille):

Hello! Could anyone help me?:) Write a formula of hydrocarbon, if you know that there is 92.3% of C and 7.7 of H. Also you know, that 1litre of this hydrocarbon (at under normal conditions) weighs 1.16g. Give me some help, please:)

OpenStudy (kamille):

And sorry for my poor English!

OpenStudy (anonymous):

find the empirical formula by dividing the percents by the molar masses of the individual elements first. 92.3/12.01 = X 7.7/1.008= Y

OpenStudy (kamille):

Thanks:)

OpenStudy (anonymous):

then divide the quotient that you get from that X and Y by the lesser of the two. ex. 7.686/ 7.63 and 7.63/7.63 Both quotients = 1. therefor your empirical formula is CH. from here you will find the molecular formula.

OpenStudy (kamille):

How can I find molecural formula? i still don't get it :\

OpenStudy (anonymous):

if the empirical formula is CH, then something is off. for hydrocarbons full saturation is noted by the formula \[C_nH_2n+2\] where n = the number of carbon atoms. the least saturated of hydrocarbons are the alkynes \[C_n H_2n-2\] C ,H enough for the empirical of CH to not make sense.

OpenStudy (anonymous):

err last line should be C<H